OpenStudy (anonymous):

Can somebody explain to me how to solve this equation? I believe that I know the answer, I just don't know how to explain it completely. u+2=4

OpenStudy (anonymous):

Move the 2 to the other side? How would you do it?

OpenStudy (shamil98):

You want to get U by itself.

OpenStudy (shamil98):

so you undo what was done to U.

OpenStudy (anonymous):

Yes you'd want to isolate u.

OpenStudy (shamil98):

2 is added to u , so you subtract 2 from both sides . u + 2 = 4 -2 -2 u = 2

OpenStudy (anonymous):

Okay, that's what I did. I just subtracted two from 4.

OpenStudy (anonymous):

What you do to one side, you must do to the other, as shamil98 showed.

OpenStudy (anonymous):

Thanks, I just forgot how to do these equations for a minute. So I guess I did it without even knowing that I did it. lol

OpenStudy (shamil98):

Btw, click on close your question, please and thank you.

OpenStudy (haseeb96):

u +2 =4 (u) keep as it is and shift +2 to the other side u=4-2 u=2 this is the method of solving it

OpenStudy (anonymous):

No problem.